# This work was supported in part by Hong Kong Innovation and Technology Commission Funding Administrative System II (ITF Ref. No. GHP/110/20GD) and the National Natural Science Foundation of China (62192714, U21B2006).(Corresponding author: Hongwei Liu; Chang Gao.)

### v2

@@ 3.2 Multi-feature Link Prediction Network
Given the challenges in deriving an optimal detector that effectively integrates multi-dimensional information, we address the problem from a data-driven perspective. By cleverly designing networks and the cost function, we achieve a near-optimal fusion detector that comprehensively utilizes diverse information on different attributes. The architecture of MFLPN is illustrated in Fig. 3. After constructing observation association graphs, NFEN first performs feature extraction on the multi-dimensional information of plots. Next, STEF-GAN is employed to extract spatio-temporal coupling features between plots, enhancing the fusion of multi-frame information. Finally, OAJN is utilized to conduct the link prediction task, yielding confidence for edges derive from target associations, and subsequently obtaining the confidence of candidate tracks originating from targets. Based on the track confidence, the second threshold $\gamma_{2}$ is employed to make detection for candidate tracks, with those exceeding $\gamma_{2}$ regarded as potential target tracks.
@@
Figure 3: The architecture of Multi-feature Link Prediction Network.
1) NFEN Layer: Information such as Doppler and echo structure can be obtained in the coherent radar system. By fully integrating these information, the algorithm performance should be enhanced. The plot information we use is as follows:
@@
where $d_{k,n}$ represents the Doppler channel in which the observation is detected, $s_{k,n}$ represents SNR, $\boldsymbol{i}_{k,n}$ represents temporal information, and $\boldsymbol{a}_{k,n}$ represents the range-Doppler map. The variation of the Doppler channel $d_{k,n}$ where the target plot is located across multiple frames reflects the target kinematic characteristics. For example, when a target moves at a constant speed, its Doppler channel remains stable or varies approximately linearly. The signal-to-noise ratio $s_{k,n}$ not only reflects the strength of the target signal but also indicates the magnitude of measurement errors. Therefore, the information held by SNR is valuable and warrants thorough exploration. The temporal information represents the chronological order of the observations in the track, serving as important temporal scale information. In this paper, the frame $n$ is converted to binary encoding as a temporal feature, denoted as $\boldsymbol{i}_{k,n}$. As shown in Fig. 4, the signal structures of target observations across multiple frames exhibit a certain similarity, which is an important feature for distinguishing targets and false alarms. After coherent integration, the corresponding range-Doppler maps which contain signal structure information for specific detection units can be obtained. We leverage range-Doppler maps to further extract characteristics of targets in the signal domain, thereby enhancing the separability between targets and false alarms.
The NFEN aims at performing feature extraction on $\boldsymbol{h}_{k,n}=\left(d_{k,n},s_{k,n},\boldsymbol{i}_{k,n},\boldsymbol{a}_{k,n}\right)^{\top}$, which in turn makes the whole network better at fusing multi-dimensional information. It contains three fully connected neural networks and one convolutional neural network. The fully connected layer is usually used for feature enhancement and transformation of dimensions. The convolutional structure not only reduces the number of parameters through parameter sharing, but also improves the learning ability of NFEN for local similarity information, which in turn leads to more effective extracting of plot features.
@@
$=\phi_{D}\left(\mathbf{W}_{d}^{\left(N_{d}\right)}\cdots\phi_{D}\left(\mathbf{W}_{d}^{(1)}d_{k,n}\right)\right)$
@@
$=\phi_{D}\left(\mathbf{W}_{s}^{\left(N_{s}\right)}\cdots\phi_{D}\left(\mathbf{W}_{s}^{(1)}s_{k,n}\right)\right)$
@@
where $D_{\cdot}^{\cdot}$ denotes the transformations between layers, $\circ$ denotes the composition of functions, $N_{d},N_{s},N_{\boldsymbol{i}}$ denote the number of fully connected layers, $\mathbf{W}_{d}^{\cdot},\mathbf{W}_{s}^{\cdot},\mathbf{W}_{\boldsymbol{i}}^{\cdot}$ denote the learnable parameter matrixes and $\phi_{D}$ denotes the nonlinear activation function in NEFN, which enhances the ability to learn nonlinear mappings. The transformation from $\boldsymbol{a}_{k,n}$ to its feature vector $\boldsymbol{f}_{k,n}^{(\boldsymbol{a})}$ is completed using a convolutional neural network:
@@
where $N_{\boldsymbol{a}}$ denotes the number of convolutional layers, * denotes the circular convolution operation and $\mathbf{W}_{\boldsymbol{a}}^{\cdot}$ denotes the parameters in the convolution kernels. Finally, the concatenation of $\boldsymbol{f}_{k,n}^{(d)},\boldsymbol{f}_{k,n}^{(s)},f_{k,n}^{(\boldsymbol{i})},f_{k,n}^{(\boldsymbol{a})}$ yields the multi-dimensional fused feature of the plot:
@@
where the fused feature $\boldsymbol{f}_{k,n}$ is used as the input node feature of the STEF-GAN layer in observation association graphs. To enhance readability, we reformulate the input node features based on the temporal order of the observations across multiple frames. Specifically, if there are $D_{l}$ observations in the $l$-th frame across $L$ frames, $l=1,2,\ldots,L$, the input node features are rewritten as $\boldsymbol{x}_{1}^{(0)},\boldsymbol{x}_{2}^{(0)},\ldots,\boldsymbol{x}_{\mathcal{L}}^{(0)}$, where $\mathcal{L}=\sum_{l=1}^{L}D_{l}$.
2) STEF-GAN Layer: Most existing Graph Neural Networks can be abstractly summarized into a universal framework known as Message Passing Neural Networks (MPNNs)[38]. MPNNs focus on a specific message passing mechanism, allowing for the transfer of node information by defining a message passing function, enabling each node to aggregate information from its neighbors and update its state accordingly. The design of contemporary GNNs mostly adheres to this message passing concept.
The observation association graphs constructed from data received by sensors at different times are distinct, indicating that the problem we address is an inductive task. Introducing an attention mechanism allows the network to focus more on globally relevant information that aids decision-making, enabling efficient resource utilization and enhancing the network’s generalization capability, particularly performing well in inductive tasks[39],[40]. Furthermore, incorporating edge features facilitates a more thorough exploitation of the spatio-temporal coupling characteristics among multi-frame observations, thereby improving the network performance. Based on these considerations, the message passing formula of the $k$-th layer in STEF-GAN is as follows:
$\boldsymbol{x}_{i}^{(k+1)}=$ $\hskip 2.84544pt\phi_{G}\left(\mathbf{W}_{G}^{(k)}\boldsymbol{x}_{i}^{(k)}\underset{j\in\mathcal{N}(i)}{\oplus}\right.$ (9)
$\left.{\color[rgb]{1,1,1}\underset{j\in\mathcal{N}(i)}{\oplus}}a_{j,i}^{(k)}\beta_{G}^{(k)}\left(\mathbf{W}_{G}^{(k)}\boldsymbol{x}_{j}^{(k)}\|\hskip 1.42271pt\mathbf{L}_{E}^{(k)}\boldsymbol{e}_{j,i}\right)\right)$
$a_{j,i}^{(k)}=$ $\hskip 2.84544pt\frac{M_{j,i}^{(k)}}{\sum_{m\in\mathcal{N}(i)}M_{m,i}^{(k)}}$
$M_{m,i}^{(k)}=$ $\hskip 2.84544pt\exp\left(\operatorname{LeakyRelu}\left(\boldsymbol{h}_{G}^{(k)}\left(\mathbf{W}_{G}^{(k)}\boldsymbol{x}_{i}^{(k)}\|\right.\right.\right.$
$\left.\left.\left.\mathbf{W}_{G}^{(k)}\boldsymbol{x}_{m}^{(k)}\|\hskip 1.42271pt\mathbf{W}_{E}^{(k)}\boldsymbol{e}_{m,i}\right)\right)\right)\text{.}$
@@
$H$ is the number of heads with $\boldsymbol{x}_{i}^{(k+1,h)}$ representing the aggregated feature of node $i$ in the $h$-th head, which independently executes the transformation shown in (9). It is precisely that the message passing mechanism which enables STEF-GAN to integrate multi-frame spatial-temporal information and reveal the relationship between observations, thereby enhancing the separability of targets and false alarms.
Figure 4: Illustration of the signal structural difference between targets and false alarms.
The time, range, azimuth and radial velocity measurements of target plots exhibit a certain coupling relationship across frames, reflecting the target physical kinematics. However, this coupling relationship is often complex. Directly using $t_{k,n},r_{k,n},\theta_{k,n},v_{k,n}$ as the node feature input to the network will increase both the training difficulty and computational burden. To address this issue, we pre-construct the mapping relationships of observation measurements across multiple frames according to the target kinematics and treat the relationships as edge features. This method effectively extracts the spatio-temporal coupling features among multi-frame observations while ensuring computational efficiency and maintaining good performance.
In practice, target motion is typically based on a certain physical mechanism. For example, considering a target moving at a constant speed, it is possible to use the position measurements in two frames to estimate the target velocity. By projecting this velocity onto the sensor line-of-sight directions in each frame, the estimation of target radial velocities in both frames can be derived. If there are no measurement errors, the estimated velocities should be equal to the velocity measurements. Considering the presence of errors, there will be a certain deviation between the estimations and the measurements, which can be used to determine whether two observations correspond to the same target. Therefore we incorporate the deviation as the edge feature in STEF-GAN to learn the confidence that two observations originate from the same target.
@@
where $v_{u}$ denotes the range of the unambiguous radial velocity while $m_{1}$ and $m_{2}$ denote the ambiguous numbers. The target moving velocity can be estimated based on the position measurements, and by projecting this velocity along the sensor line-of-sight, the target estimated radial velocities can be obtained:
$\hat{v}_{1}=\frac{r_{2}\cos\left(\theta_{2}-\theta_{1}\right)-r_{1}}{t_{2}-t_{1}}$ (12)
$\hat{v}_{2}=\frac{r_{2}-r_{1}\cos\left(\theta_{2}-\theta_{1}\right)}{t_{2}-t_{1}}$
@@
$F_{MC,1}=\left|v_{1}+m_{1}v_{u}-\hat{v}_{1}\right|$ (13)
$F_{MC,2}=\left|v_{2}+m_{2}v_{u}-\hat{v}_{2}\right|$
$m_{1}=\underset{m\in\mathbb{Z}}{\arg\min}\left|v_{1}+m_{1}v_{u}-\hat{v}_{1}\right|$
@@
Since $F_{MC,1}$ and $F_{MC,2}$ are features derived from the information of two nodes, they can serve as edge features in observation association graphs, denoted as $\boldsymbol{e}_{1,2}=\left[F_{MC,1},F_{MC,2}\right]$.
In summary, after obtaining the input features for nodes and edges in observation association graphs, the STEF-GAN leverages multi-layer message passing to extract information from multiple frames:
$\boldsymbol{X}^{\left(N_{G}\right)}=\left(G^{\left(N_{G}\right)}\circ\ldots\circ G^{(2)}\circ G^{(1)}\right)(\boldsymbol{x}^{(0)},\boldsymbol{e})\text{,}$ (14)
where $G^{\cdot}$ denotes the transformations between layers and $N_{G}$ is the number of layers. $\boldsymbol{x}^{(0)}$ represents the node feature matrix input to STEF-GAN, which is also the plot fused feature matrix $\boldsymbol{f}$ output from NEFN and $\boldsymbol{e}$ represents the edge feature matrix.
3) OAJN Layer: After thoroughly extracting multi-dimensional features of plots and spatio-temporal coupling features among multi-frame plots, we employ OAJN to achieve intelligent and integrated track search and track detection process within observation association graphs. Compared to traditional methods that utilize energy accumulation compared with a given threshold for track detection[18], this approach excels in learning the mapping from measurement information to whether plot associations stem from the target, thus providing a better assessment of the track quality. The proposed method first accomplishes a link prediction task, classifying edges to obtain the confidence that the plot associations originate from the target. Subsequently, it determines the confidence that the candidate tracks are derived from the target and conducts track detection based on a second detection threshold $\gamma_{2}$.
Considering that $\boldsymbol{e}_{j,i}=\left[F_{MC,j},F_{MC,i}\right]$ directly reflects the coupling relationship between position and Doppler measurements, we reintegrate edge features into OAJN for more effective utilization. Additionally, the difference in Doppler channels of the target is an important feature for target motion characteristics. For instance, the difference in Doppler channels of a target moving at a constant speed across different frames is typically small. As a consequence, we also incorporate the Doppler channel difference $d_{j,i}=\left|d_{j}-d_{i}\right|$ between node $i$ and node $j$ into OAJN to enhance the learning of link prediction. To facilitate the subsequent distinguishing between target tracks and false tracks based on the second threshold, we select the softmax function as the nonlinear activation when decoding. In the observation association graph, edges are categorized into three types: the edge between false alarms, the edge between a false alarm and a target, and the edge between targets. Their corresponding ground truths are denoted as $y_{j,i}^{(1)}$, $y_{j,i}^{(2)}$, $y_{j,i}^{(3)}$. For any two connected nodes $i$ and $j$, the classification of their edge is expressed as follows:
@@
$=\operatorname{softmax}\left(\mathbf{W}_{J}^{\left(N_{J}\right)}\cdots\phi_{J}\left(\mathbf{W}_{J}^{(1)}\boldsymbol{F}_{j,i}\right)\right)$
$\boldsymbol{F}_{j,i}$ $=\left[\boldsymbol{x}_{i}^{\left(N_{G}\right)};\boldsymbol{x}_{j}^{\left(N_{G}\right)};\boldsymbol{e}_{j,i};d_{j,i}\right]\text{,}$
where $\hat{y}_{j,i}^{(c)},c=1,2,3$ represents the classification score for edges, indicating the confidence that the edges originate from target associations. $\boldsymbol{x}_{j}^{\left(N_{G}\right)}$ and $\boldsymbol{x}_{i}^{\left(N_{G}\right)}$ are the output node features of STEF-GAN, while $\boldsymbol{F}_{j,i}$ is the concatenated feature. $\mathbf{W}_{J}^{\cdot}$ denotes the learnable parameter matrix in OAJN and $\phi_{J}$ is the nonlinear activation function. The categorical cross-entropy loss function is employed in this work, which not only alleviates the gradient vanishing problem caused by deep networks but is also equivalent to obtaining a nonlinear mapping from multi-dimensional information in raw data to target detection through maximum likelihood estimation[43]. The categorical cross-entropy loss function used for integrated training is formulated as follows:
@@
The confidence $S$ that a candidate track originates from a target is defined as follows:
$S$ $=\frac{1}{N_{E}}\sum_{e=1}^{N_{E}}s_{e}+\lambda N_{V}$ (17)
$s_{e}$ $=\alpha_{1}\cdot\hat{y}_{e}^{(1)}+\alpha_{2}\cdot\hat{y}_{e}^{(2)}+\alpha_{3}\cdot\hat{y}_{e}^{(3)}\text{,}$
where $N_{E}$ represents the edge number of the track, $N_{V}$ represents the track length, $\hat{y}_{e}^{(c)},c=1,2,3$ represents the classification score of the $e$-th edge, $\alpha_{c}$ and $\lambda$ are both hyperparameters. For the configuration of hyperparameters, we offers some empirical recommendations. For a target track, the most desirable edges are those between targets, while the least desirable are those between false alarms. Therefore, the weight for the classification score of edges between targets can be set to $\alpha_{3}=1$, while the weight for edges between false alarms can be set to $\alpha_{1}=0$. As to the weight for edges between targets and false alarms, it can be set to a small positive value, such as $\alpha_{2}=0.2$. On one hand, we aim to preserve as many edges between targets as possible, while also striving to suppress edges between targets and false alarms. Thus, the weight should not be too large. On the other hand, to differentiate from edges between false alarms, and considering the presence of target plots, the weight should preferably be greater than $\alpha_{1}$. Additionally, without affecting subsequent target tracking, associations between targets and false alarms can help to improve track continuity. For example, in scenarios where a weak target is not detected in the current frame, if an association occurs between the historical target track and a false alarm without negatively impacting subsequent target tracking, it can actually improve track continuity.
The weight $\lambda$ of track length should be balanced with the weights of edges. From one perspective, it should not be set too high, as this may place excessive emphasis on track length, leading to an increase in false alarms within the track. From another perspective, while ensuring that the confidence score for edges originating from target associations is high, longer tracks should be chosen to enhance track continuity. After obtaining the track confidence, a second detection threshold $\gamma_{2}$ is applied for track detection. Tracks that exceed $\gamma_{2}$ are classified as potential target tracks. Finally, following track pruning, the confirmed tracks are generated and output.

### v3

@@ 3.2 Multi-feature link prediction network
Considering the challenges in deriving a detector that integrates multi-dimensional information, we address the problem from a data-driven perspective [40]. By cleverly designing networks and the cost function, we perform fusion detection by comprehensively utilizing diverse information. The architecture of MFLPN is illustrated in Fig. 3. After constructing observation association graphs, the node feature embedding network (NFEN) is developed to extract plot features such as Doppler information and signal structures by leveraging the strong fitting capabilities of deep neural networks (DNNs) for nonlinear mapping [41]. Subsequently, we design the graph attention network with spatio-temporal coupling encoded in edge features (STEF-GAT) based on GNNs. By incorporating an attention mechanism and constructing edge features during message passing, the network adequately extracts relational information that embodies physical motion patterns among target observations. Then, we develop the observation association judgment network (OAJN) to conduct the link prediction task in graphs, yielding confidence for edges deriving from target associations. And the confidence score for each candidate track originating from targets can be obtained. Finally, based on the track confidence, the final detection threshold $\gamma_{2}$ is employed to make detection for candidate tracks, with those exceeding $\gamma_{2}$ declared as potential tracks.
@@
Figure 3: The architecture of the multi-feature link prediction network.
1) NFEN: Information such as the Doppler and signal structure can be obtained in a coherent system, while fully integrating the information of the target distributed in the echo should be helpful to detect the target from the background [42],[43],[44]. Based on (1), all the observation information we use is as follows:
@@
where $d_{k,n}$ represents the Doppler channel in which the observation is detected, $s_{k,n}$ represents SNR, $\boldsymbol{i}_{k,n}$ represents temporal information, and $\boldsymbol{a}_{k,n}$ represents the range-Doppler map. The variation of the Doppler channel $d_{k,n}$ where the target observation is located across multiple frames reflects the target kinematic characteristics. For example, when a target moves at a constant speed, its Doppler channel remains stable or varies approximately linearly. $s_{k,n}$ not only reflects the strength of the target signal but also indicates the magnitude of measurement errors. Therefore, the information contained in SNR is valuable and warrants thorough exploration. The temporal information represents the chronological order of the observations in the track, serving as important temporal scale information. In this paper, the frame $n$ is converted to binary encoding as a temporal feature, denoted as $\boldsymbol{i}_{k,n}$. As shown in Fig. 4, the signal structures of target observations across multiple frames exhibit a certain similarity, which is an important feature for distinguishing targets and false alarms. After coherent integration, the corresponding range-Doppler maps which contain signal structure information for specific detection cells can be obtained. We leverage range-Doppler maps to further extract characteristics of targets in the signal domain, thereby enhancing the separability between targets and false alarms.
The NFEN aims to perform node feature extraction on $\boldsymbol{h}_{k,n}=\left(d_{k,n},s_{k,n},\boldsymbol{i}_{k,n},\boldsymbol{a}_{k,n}\right)^{\top}$, which enables the whole network to better fuse multi-dimensional information. It contains three fully connected neural networks and one convolutional neural network. The fully connected layer is usually used for transformation of feature dimensions. The convolutional structure not only reduces the number of parameters through parameter sharing but also improves the learning ability of NFEN for local similarity information, which in turn leads to more effective extraction of plot features.
@@
$=\phi_{D}\left(\mathbf{W}_{d}^{\left(N_{d}\right)}\cdots\phi_{D}\left(\mathbf{W}_{d}^{(1)}d_{k,n}\right)\right)\text{,}$
@@
$=\phi_{D}\left(\mathbf{W}_{s}^{\left(N_{s}\right)}\cdots\phi_{D}\left(\mathbf{W}_{s}^{(1)}s_{k,n}\right)\right)\text{,}$
@@
where $D_{\cdot}^{\cdot}$ denotes the transformations between layers, $\circ$ denotes the composition of functions, $N_{d},N_{s},N_{\boldsymbol{i}}$ denote the number of fully connected layers, $\mathbf{W}_{d}^{\cdot},\mathbf{W}_{s}^{\cdot},\mathbf{W}_{\boldsymbol{i}}^{\cdot}$ denote the learnable parameter matrixes and $\phi_{D}$ denotes the nonlinear activation function in NFEN, which enhances the ability to learn nonlinear mappings. The transformation from $\boldsymbol{a}_{k,n}$ to its feature vector $\boldsymbol{f}_{k,n}^{(\boldsymbol{a})}$ is completed using a convolutional neural network:
@@
where $N_{\boldsymbol{a}}$ denotes the number of convolutional layers, * denotes the circular convolution operation and $\mathbf{W}_{\boldsymbol{a}}^{\cdot}$ denotes the parameters in the convolution kernels. Finally, the concatenation of $\boldsymbol{f}_{k,n}^{(d)},\boldsymbol{f}_{k,n}^{(s)},\boldsymbol{f}_{k,n}^{(\boldsymbol{i})},\boldsymbol{f}_{k,n}^{(\boldsymbol{a})}$ yields the multi-dimensional fused feature of the observation:
@@
where the fused feature $\boldsymbol{f}_{k,n}$ is used as the input node feature of the STEF-GAT in observation association graphs. To enhance readability, we reformulate the input node features based on the temporal order of the observations across multiple frames. Specifically, if there are $D_{l}$ observations in the $l$-th frame across $L$ frames, $l=1,2,\ldots,L$, the input node features are rewritten as $\boldsymbol{x}_{1}^{(0)},\boldsymbol{x}_{2}^{(0)},\ldots,\boldsymbol{x}_{\mathcal{L}}^{(0)}$, where $\mathcal{L}=\sum_{l=1}^{L}D_{l}$.
2) STEF-GAT: Most existing GNNs can be abstractly summarized into a universal framework known as message passing neural networks (MPNNs) [45]. MPNNs focus on a specific message passing mechanism, allowing for the transfer of node information by defining a message passing function. Therefore, MPNNs enable each node to aggregate information from its neighbors and update its state accordingly. The design of contemporary GNNs mostly adheres to this message passing concept.
The observation association graphs constructed from data received by sensors at different times are distinct, indicating that the problem we address is an inductive task. Introducing an attention mechanism allows the network to focus more on relevant information that is crucial for decision-making, enabling efficient resource utilization and enhancing the generalization capability of the network, particularly performing well in inductive tasks [46],[47]. Furthermore, incorporating edge features facilitates a more thorough exploitation of the spatio-temporal coupling information among multi-frame observations, thereby improving the network performance. Based on these considerations, the message passing formula of the $k$-th layer in STEF-GAT is designed as follows:
$\boldsymbol{x}_{i}^{(k+1)}=$ $\hskip 2.84544pt\phi_{G}\left(a_{i,i}^{(k)}\mathbf{W}_{G}^{(k)}\boldsymbol{x}_{i}^{(k)}\underset{j\in\mathcal{N}(i)}{\oplus}\right.$ (10)
$\left.{\color[rgb]{1,1,1}\underset{j\in\mathcal{N}(i)}{\oplus}}a_{j,i}^{(k)}\beta_{G}^{(k)}\left(\mathbf{W}_{G}^{(k)}\boldsymbol{x}_{j}^{(k)}\|\hskip 1.42271pt\mathbf{L}_{E}^{(k)}\boldsymbol{e}_{j,i}\right)\right)\text{,}$
$a_{j,i}^{(k)}=$ $\hskip 2.84544pt\frac{\exp\left(M_{j,i}^{(k)}\right)}{\sum_{m\in\mathcal{N}(i)\cup\{i\}}\exp\left(M_{m,i}^{(k)}\right)}\hskip 2.84544pt\text{,}$
$M_{m,i}^{(k)}=$ $\hskip 2.84544pt\operatorname{LeakyRelu}\left(\boldsymbol{a}_{G}^{(k)}\left(\mathbf{W}_{G}^{(k)}\boldsymbol{x}_{i}^{(k)}\|\right.\right.$
@@
[…]
$H$ represents the number of heads. $\boldsymbol{x}_{i}^{(k+1,h)}$ represents the feature of node $i$ in the $h$-th head, which independently executes the transformation shown in (10). It is precisely that the message passing mechanism which enables STEF-GAT to integrate multi-frame spatial-temporal information, reveals the relationships among multi-frame observations, thereby enhancing the separability of targets and false alarms.
Figure 4: Illustration of the signal structure difference between targets and false alarms.
The time, range, azimuth and radial velocity measurements of target observations exhibit a certain coupling relationship across frames, reflecting the target physical kinematics. However, this coupling relationship is often complex. Directly using $t_{k,n},r_{k,n},\theta_{k,n},v_{k,n}$ as the node feature input to the network will require deeper network layers, increasing both the training difficulty and computational burden. To address this issue, we pre-construct the mapping relationships of observation measurements across multiple frames and treat the relationships as edge features. This method effectively extracts the spatio-temporal coupling information among multi-frame observations while reducing computational load and maintaining good performance. In practice, target motion is typically based on a certain physical mechanism. For example, considering a target moving at a constant speed, it is possible to use the position measurements in two frames to estimate the target velocity. By projecting this velocity onto the sensor line-of-sight directions in each frame, the estimation of target radial velocities in both frames can be derived. If there are no measurement errors, the estimated velocities should be equal to the velocity measurements. Considering the presence of errors, there will be a certain deviation between the estimations and the measurements, which can be used to determine whether two observations correspond to the same target. Therefore, we incorporate the deviation as the edge feature in STEF-GAT to learn the confidence that two observations originate from the same target.
@@
where $v_{u}$ denotes the range of the unambiguous radial velocity, while $m_{1}$ and $m_{2}$ denote the ambiguous numbers. The target estimated radial velocities can be obtained based on the position measurements:
$\hat{v}_{1}=\frac{r_{2}\cos\left(\theta_{2}-\theta_{1}\right)-r_{1}}{t_{2}-t_{1}}\text{,}$ (13)
$\hat{v}_{2}=\frac{r_{2}-r_{1}\cos\left(\theta_{2}-\theta_{1}\right)}{t_{2}-t_{1}}\text{.}$
@@
$F_{E,1}$ $=\left|v_{1}+m_{1}v_{u}-\hat{v}_{1}\right|\text{,}$ (14)
$F_{E,2}$ $=\left|v_{2}+m_{2}v_{u}-\hat{v}_{2}\right|\text{,}$
$m_{1}$ $=\underset{m\in\mathbb{Z}}{\arg\min}\left|v_{1}+m_{1}v_{u}-\hat{v}_{1}\right|\text{,}$
@@
Since $F_{E,1}$ and $F_{E,2}$ are features derived from the information of two nodes, they can serve as edge features in observation association graphs, denoted as $\boldsymbol{e}_{1,2}=\left[F_{E,1},F_{E,2}\right]$.
After obtaining the input features of nodes and edges for STEF-GAT, it leverages multi-layer message passing to extract information from multiple frames:
[…]
where $G^{\cdot}$ denotes the transformations between layers and $N_{G}$ is the number of layers. $\boldsymbol{x}^{(0)}$ represents the node feature matrix input to STEF-GAT, which is also the fused feature matrix $\boldsymbol{f}$ output from NFEN. And $\boldsymbol{e}$ represents the edge feature matrix.
3) OAJN: After thoroughly extracting multi-dimensional information of plots and spatio-temporal coupling information among multi-frame plots, we employ OAJN to achieve intelligent and integrated track search and track detection process within observation association graphs. Compared to traditional methods that utilize energy accumulation with a given threshold for track detection [18], this approach excels in learning the mapping from measurement information to whether observation associations stem from the target, thus providing a better assessment of the track quality. First, OAJN accomplishes a link prediction task, classifying edges to obtain the confidence that the observation associations originate from the target. Subsequently, the confidence that the candidate tracks are derived from the target can be obtained and the track detection is performed using the final detection threshold $\gamma_{2}$.
Figure 5: Message passing in STEF-GAT.
Considering that $\boldsymbol{e}_{j,i}=\left[F_{E,j},F_{E,i}\right]$ directly reflects the coupling relationship between position and Doppler measurements, we reintegrate edge features into OAJN for more effective utilization. Additionally, the difference in Doppler channels of the target is an important feature for target kinematic characteristics. For instance, the difference in Doppler channels of a target moving at a constant speed across different frames is typically small. As a consequence, we also incorporate the Doppler channel difference $d_{j,i}=\left|d_{j}-d_{i}\right|$ between node $i$ and node $j$ into OAJN to enhance the learning of link prediction. To facilitate the subsequent distinguishing between target tracks and false tracks based on the final detection threshold, we select the softmax function as the nonlinear activation when decoding. In the observation association graph, edges are categorized into three types: the edge between false alarms, the edge between a false alarm and a target, and the edge between targets. Their corresponding desired labels are denoted as $y_{j,i}^{(1)}$, $y_{j,i}^{(2)}$, $y_{j,i}^{(3)}$. For any two connected nodes $i$ and $j$, the classification of their edge is expressed as follows:
@@
$=\operatorname{softmax}\left(\mathbf{W}_{J}^{\left(N_{J}\right)}\cdots\phi_{J}\left(\mathbf{W}_{J}^{(1)}\boldsymbol{F}_{j,i}\right)\right)\text{,}$
$\boldsymbol{F}_{j,i}$ $=\left(\boldsymbol{x}_{i}^{\left(N_{G}\right)}\|\hskip 1.9919pt\boldsymbol{x}_{j}^{\left(N_{G}\right)}\|\hskip 1.9919pt\mathbf{W}_{ed}\left(\boldsymbol{e}_{j,i}\|d_{j,i}\right)\right)\text{,}$
[…]
@@
4) Track detection: The confidence $S$ that a candidate track originates from a target is defined as follows:
$S$ $=\frac{1}{N_{E}}\sum_{e=1}^{N_{E}}\rho_{e}+\lambda N_{V}\text{,}$ (18)
$\rho_{e}$ $=\alpha_{1}\cdot\hat{y}_{e}^{(1)}+\alpha_{2}\cdot\hat{y}_{e}^{(2)}+\alpha_{3}\cdot\hat{y}_{e}^{(3)}\text{,}$
where $N_{E}$ represents the number of edges in the track, $N_{V}$ represents the track length, $\rho_{e}$ represents the confidence for the edge deriving from the target association, $\hat{y}_{e}^{(c)},c=1,2,3$ represents the classification score of the $e$-th edge, $\alpha_{c}$ and $\lambda$ are both hyperparameters. For the configuration of hyperparameters, we offer some empirical recommendations. For a target track, the most desirable edges are those between targets, while the least desirable are those between false alarms. Therefore, the weight for the classification score of edges between targets can be set to $\alpha_{3}=1$, while the weight for edges between false alarms can be set to $\alpha_{1}=0$. Regarding the weight for edges between targets and false alarms, it can be set to a small positive value, such as $\alpha_{2}=0.2$. On one hand, we aim to preserve as many edges between targets as possible, while also striving to suppress edges between targets and false alarms. Thus, the weight should not be too large. On the other hand, to differentiate from edges between false alarms, and considering the presence of target observations, the weight should preferably be larger than $\alpha_{1}$. Additionally, without affecting subsequent target tracking, associations between targets and false alarms can help to improve track continuity. For example, in scenarios where a weak target is not detected in the current frame, if an association occurs between the historical target track and a false alarm without negatively impacting subsequent target tracking, it can actually improve track continuity.
The weight $\lambda$ of track length should be balanced with the weights of edges. From one perspective, it should not be set too high, as this may place excessive emphasis on track length, leading to an increase in false alarms within the track. From another perspective, while ensuring that the confidence score for edges originating from target associations is high, longer tracks should be chosen to enhance track continuity. After obtaining the track confidence, the final detection threshold $\gamma_{2}$ is applied for track detection. Candidate tracks with the confidence score exceeding $\gamma_{2}$ are considered as potential target tracks. Finally, following track pruning, the confirmed tracks are generated and output.